OpenStudy (mackenzie2013):

The function y = x + 7 is graphed in the coordinate plane. Which point will not be on the line? (0, 7) (2, 14) (9, 16) (12, 19)

jimthompson5910 (jim_thompson5910):

plug in each answer choice if you get a true equation, then that point is on the line

jimthompson5910 (jim_thompson5910):

if you get a false equation, the point is not on the line

OpenStudy (anonymous):

Plug in x values for x and solve for y tell me what you get

OpenStudy (mackenzie2013):

2,14?

jimthompson5910 (jim_thompson5910):

correct again

jimthompson5910 (jim_thompson5910):

y = x+7 14 = 2+7 .. plug in (2,14) 14 = 9 ... false ---> (2,14) is not on the line y =x+7